for each data set a. determine a frequency distribution b. obtain a relative frequency distribuation c.draw a pie chart d. Consruct a bar graph


1.C A B B A


2.D D B D D A A C A A
Answer #1

1.

Data Frequency Relative Frequency
A 2 2/5 * 100 = 40%
B 2 2/5 * 100 = 40%
C 1 1/5 * 100 = 20%
Total 5 100%

The pie chart and bar graph is shown below.

2.

Data Frequency Relative Frequency
A 4 4/10 * 100 = 40%
B 1 1/10 * 100 = 10%
C 1 1/10 * 100 = 10%
D 4 4/10 * 100 = 40%
Total 10 100%

The pie chart and bar graph is shown below.